Latasha can read 40 pages of astronomy in an hour. She can also read 30 pages of sociology in an hour. She spends 4 hours per day studying. Use the blue line (circle symbol) to draw Latasha's production possibilities frontier (PPF) for reading astronomy and sociology. GRAPH HERE Latasha's opportunity cost of reading 90 pages of sociology is pages of astronomy.
Answer #1

Ans:

Latasha's production possibilities frontier (PPF) for reading astronomy and sociology is shown in the following figure.

Latasha can read 40 pages of astronomy in an hour.

Latasha can read total pages of astronomy in 4 hour = 40 * 4 = 160 pages

Latasha can read 30 pages of sociology in an hour.

Latasha can read total pages of sociology in 4 hour = 30 * 4 = 120 pages

Ans: Latasha's opportunity cost of reading 90 pages of sociology is 120 pages of astronomy.

Explanation:

Opportunity cost of reading 1 page of sociology = 160 / 120 = 1.33 pages of astronomy.

Opportunity cost of reading 90 pages of sociology = 1.33 * 90 = 120 pages of astronomy.